OpenStudy (anonymous):

anyone know -(z+6)+(4z+1)=-4(z+1)

OpenStudy (anonymous):

z=-1

OpenStudy (unklerhaukus):

\[-(z+6)+(4z+1)=-4(z+1)\]\[-z-6+4z+1=-4z-4\] \[-z+4z+4z=-4+6-1\]

OpenStudy (anonymous):

Wait never mind my answer is wrong

OpenStudy (anonymous):

-z-6+4z+1=-4z-4 7z=1 z=(1/7)

OpenStudy (anonymous):

thank u so much